I need to find out how many gumballs are in a gumball machine. If the diameter of the sphere is 4cm and if a square of 9 gumballs is 1/2cm. I am clueless. Please help.
Answer by checkley71(8403)

V=4/3PIR^3
V=4/3*3.14*2^3
V=9.42/3*8
V=25.12 ANSWER FOR THE MACHINE VOLUME.
9/0.5=X/25.12 NOW CROSS MULTIPLY.
.5X=226.08
X=226.08/.5
X=452.16 ANSWER FOR THE NUMBER OF BALLS.
----------------------------------------------
THIS ANSWER ASSUMES THE BALLS ALL STACK UP IN A SQUARE. HOWEVER: THAT ISN'T THE WAY THEY LIE IN THE MACHINE SO THE ACTUAL TOTAL NUMBER MAY BE SLIGHTLY LARGER.